# ControlMM: Controllable Masked Motion Generation

### v1

@@ Abstract
Recent advances in motion diffusion models have enabled spatially controllable text-to-motion generation. However, despite achieving acceptable control precision, these models suffer from generation speed and fidelity limitations. To address these challenges, we propose ControlMM, a novel approach incorporating spatial control signals into the generative masked motion model. ControlMM achieves real-time, high-fidelity, and high-precision controllable motion generation simultaneously. Our approach introduces two key innovations. First, we propose masked consistency modeling, which ensures high-fidelity motion generation via random masking and reconstruction, while minimizing the inconsistency between the input control signals and the extracted control signals from the generated motion. To further enhance control precision, we introduce inference-time logit editing, which manipulates the predicted conditional motion distribution so that the generated motion, sampled from the adjusted distribution, closely adheres to the input control signals. During inference, ControlMM enables parallel and iterative decoding of multiple motion tokens, allowing for high-speed motion generation. Extensive experiments show that, compared to the state of the art, ControlMM delivers superior results in motion quality, with better FID scores (0.061 vs 0.271), and higher control precision (average error 0.0091 vs 0.0108). ControlMM generates motions 20 times faster than diffusion-based methods. Additionally, ControlMM unlocks diverse applications such as any joint any frame control, body part timeline control, and obstacle avoidance. Video visualization can be found at https://exitudio.github.io/ControlMM-page/
Figure 1: ControlMM enables a wide range of applications in text-to-motion generation with high quality and precision. (a) Any Joint, Any Frame Control: spatial control signals for specific joints and frames. (b) Object Avoidance for All Joints: generates motion that avoids obstacles for any joint. (c) Body Part Timeline Control: generates motion from multiple text prompts, each corresponding to different body parts.
@@ 1 Introduction
Text-driven human motion generation has recently gained significant attention due to the semantic richness and intuitive nature of natural language descriptions. This approach has broad applications in animation, film, virtual/augmented reality (VR/AR), and robotics. While text descriptions offer a wealth of semantic guidance for motion generation, they often fall short in providing precise spatial control over specific human joints, such as the pelvis and hands. As a result, achieving natural interaction with the environment and fluid navigation through 3D space remains a challenge.
@@
To tackle this challenge, a few controllable motion generation models have been developed recently to synthesize realistic human movements that align with both text prompts and spatial control signals Shafir et al. (2023); Rempe et al. (2023); Xie et al. (2023); Wan et al. (2023). However, existing solutions face significant difficulties in generating high-fidelity motion with precise and flexible spatial control while ensuring real-time inference. In particular, current models struggle to support both sparse and dense spatial control signals simultaneously. For instance, some models excel at generating natural human movements that traverse sparse waypoints Karunratanakul et al. (2023); Rempe et al. (2023), while others are more effective at synthesizing motions that follow detailed trajectories specifying human positions at each time point Wan et al. (2023). Recent attempts to support both sparse and dense spatial inputs encounter issues with control precision; the generated motion often is not aligned well with the control conditions Xie et al. (2023). Besides unsatisfied spatial flexibility and accuracy, the quality of motion generation in controllable models remains suboptimal, as evidenced by much worse FID scores compared to models that rely solely on text inputs. Moreover, most current methods utilize motion-space diffusion models, applying diffusion processes directly to raw motion sequences. While this design facilitates the incorporation of spatial control signals, the redundancy in raw data introduces computational overhead, resulting in slower motion generation speeds.
To address these challenges, we present ControlMM, a novel approach that integrates spatial control signals into generative masked motion models that excels in high-quality and fast motion generation Pinyoanuntapong et al. (2024b); Guo et al. (2023); Pinyoanuntapong et al. (2024a). ControlMM is the first method capable of achieving real-time, high-fidelity, and high-precision controllable motion generation simultaneously. Our contributions can be summarized as follows.
  • We introduce masked consistency modeling, the first approach that incorporates spatial guidance into Masked Motion Model, which results in higher generation quality, more precise control, accelerated generation, and broader applications compared to existing methods as shown in Fig. 1.
  • We propose an inference-time logit-editing approach, which strikes the optimal balance between inference time and control precision, while enabling new control tasks, such as obstacle avoidance, in a zero-shot manner.
@@
  • We conduct extensive qualitative and quantitative evaluations on multiple tasks. As shown in Fig. 2, our model outperforms current state-of-the-art methods in motion generation quality, control precision, and speed with multiple applications i.e. joint-specific control, obstacle avoidance, body part timeline Control.
@@ 2 Related Work
Text-driven Motion Generation. Early methods for text-to-motion generation primarily focus on aligning the latent distributions of motion and language, typically by employing loss functions such as Kullback-Leibler (KL) divergence and contrastive losses. Representative works in this domain include Language2Pose Ahuja & Morency (2019), TEMOS Petrovich et al. (2022), T2M Guo et al. (2022b), MotionCLIP Tevet et al. (2022a), and DropTriple Yan et al. (2023). However, the inherent discrepancy between the distribution of text and motion often results in suboptimal generation quality when using these latent space alignment techniques.
@@
Recently, diffusion models have become a widespread choice for text-to-motion generation, operating directly in the motion space (Tevet et al., 2022b; Zhang et al., 2022; Kim et al., 2022), VAE latent space (Chen et al., 2022), or quantized space (Lou et al., 2023; Kong et al., 2023). In these works, the model gradually denoises the whole motion sequence to generate the output in the reverse diffusion process. Another line of work explores the token-based models in the human motion domain, for example, autoregressive GPTs (Guo et al., 2022a; Zhang et al., 2023a; Jiang et al., 2023; Zhong et al., 2023) and masked motion modeling (Pinyoanuntapong et al., 2024b; a; Guo et al., 2023; Hosseyni et al., 2024). These methods learn to generate discrete motion token sequences that are obtained from a pretrained motion VQVAE Esser et al. (2020); Williams et al. (2020). While GPT models usually predict the next token from history tokens, masked motion models utilize the bidirectional context to decode the masked motion tokens. By predicting multiple tokens at once, the masked modeling methods can generate motion sequences in as few as 15 steps, achieving state-of-the-art performance on generation quality and efficiency.
@@
Controllable Motion Synthesis. In addition to text prompts, synthesizing motion based on other control signals has also been a topic of interest. Example control modalities include music (Li et al., 2021b; a; Lee et al., 2019; Siyao et al., 2022; 2023; Tseng et al., 2022), interacting object (Kulkarni et al., 2024; Diller & Dai, 2024; Li et al., 2023; Cha et al., 2024), tracking sensors (Du et al., 2023), scene (Huang et al., 2023; Wang et al., 2024) programmable motion (Liu et al., 2024), style (Zhong et al., 2024), goal-reaching task (Diomataris et al., 2024), and multi-Track timeline Control (Petrovich et al., 2024). Peng et al. (2021; 2022); Xie et al. (2021); Yuan et al. (2022); Luo et al. (2023a; b); Tessler et al. (2024) incorporate physics to motion generation. To control the trajectory, PriorMDM (Shafir et al., 2023) finetunes MDM to enable control over the locations of end effectors. GMD (Karunratanakul et al., 2023) and Trace and Pace (Rempe et al., 2023) incorporates spatial control into the diffusion process by guiding the root joint location. OmniControl (Xie et al., 2023) extends the control framework to any joint, while MotionLCM (Dai et al., 2024) applies this control in the latent space, both leveraging ControlNet (Zhang et al., 2023b). DNO (Karunratanakul et al., 2024) introduces an optimization process on the diffusion noise to generate motion that minimizes a differentiable objective function. Recent approaches (Wan et al., 2023; Huang et al., 2024) model each body part separately to achieve fine-grained control but are limited to dense trajectory objectives. While effective, these methods are either limited to the specific signals, achieve imprecise control, or rely on optimizing over the diffusion model which is costly and impractical for real-time applications.
3 ControlMM
@@
The objective of ControlMM is to enable controllable text-to-motion generation based on a masked motion model that simultaneously delivers high precision, high speed, and high fidelity. In particular, given a text prompt and an additional spatial control signal, our goal is to generate a physically plausible human motion sequence that closely aligns with the textual descriptions, while following the spatial control conditions, i.e., $(x,y,z)$ positions of each human joint at each frame in the motion sequence. Towards this goal, in Section 3.1, we first introduce the background of conditional motion synthesis based on the generative masked motion model. We then describe two key components of ControlMM, including masked consistency training in Section 3.2 and inference-timelogits editing in Section 3.3. The first component aims to learn the categorical distribution of motion tokens, conditioned on spatial control during training time. The second component aims to improve control precision by optimally modifying learned motion distribution via logits editing during inference time.
Figure 3: (a) Pretrained Motion Tokenizer from a masked motion model. (b) Training phase of ControlMM, the pretrained Encoder, Decoder and Conditioned Masked Transformer are frozen, only the Motion Control Model is trained.
3.1 Preliminary: Generative Masked Motion Model
@@
Masked Motion Models consist of two stages: Motion Tokenizer and Text-conditioned Masked Transformer. The objective of the Motion Tokenizer is to learn a discrete representation of motion by quantizing the encoder’s output embedding $z$ into a codebook $\mathcal{C}$. For a given motion sequence $\mathcal{X}=[x_{1},x_{2},...,x_{T}]$, where each frame $x$ represents a 3D pose, the encoder compresses $\mathcal{X}$ into a latent embedding $z\in\mathbb{R}^{t\times d}$ with a downsampling rate of $T/t$. The embedding $z$ is quantized into codes $c\in\mathcal{C}$ from the codebook $\mathcal{C}=\{\gamma_{k}\}_{k=1}^{K}$, which contains $K$ codes. The nearest code is selected by minimizing the Euclidean distance between $z$ and the codebook entries, computed as $\hat{z_{i}}=\operatorname{argmin}_{j}\left\|\mathbf{z}-\mathbf{\mathcal{C}}_{j}\right\|_{2}^{2}$. The vector quantization loss $L_{VQ}$ is defined as:
$L_{VQ}=\|\operatorname{sg}(\mathbf{z})-\mathbf{e}\|_{2}^{2}+\beta\|\mathbf{z}-\operatorname{sg}(\mathbf{e})\|_{2}^{2},$ (1)
@@
During the second stage, the quantized motion token sequence $\mathcal{X}$ is updated with [MASK] tokens to form the corrupted motion sequence $X_{\overline{\mathbf{M}}}$. This corrupted sequence along with text embedding $W$ are fed into a text-conditioned masked transformer parameterized by $\theta$ to reconstruct input motion token sequence with reconstruction probability equal to $p_{\theta}\left(x_{i}\mid X_{\overline{\mathbf{M}}},W\right)$. The objective is to minimize the negative log-likelihood of the predicted masked tokens conditioned on text:
@@
$\mathcal{L}_{\text{mask}}=-\underset{\mathbf{X}\in\mathcal{D}}{\mathbb{E}}\left[\sum_{\forall i\in[1,L]}\log p\left(x_{i}\mid X_{\overline{\mathbf{M}}},W\right)\right].$ (2)
@@
During inference, the transformer masks out the tokens with the least confidence and predicts them in parallel in the subsequent iteration. The number of masked tokens $n_{M}$ is controlled by a masking schedule, a decaying function of the iteration $t$. Early iterations use a large masking ratio due to high uncertainty, and as the process continues, the ratio decreases as more context is available from previous predictions.
3.2 Masked Consistency Modeling
@@
ControlMM aims to generate a human motion sequence based on the text prompt (W) and spatial control signal (S). Towards this goal, we introduce a masked consistency modelling approach, which aims to learn the motion token distribution jointly conditioned on $W$ and $S$ by exploiting conditional token masking with consistency feedback.
@@
Conditioned Masked Transformer with Motion Control Model. We design a masked transformer architecture to learn the conditional motion token distribution. This is the first attempt to incorporate the ControlNet design principle (Zhang et al., 2023b) from diffusion models into generative masked models, such as BERT-like models for image, video, language, and motion generation (Devlin et al., 2019; Chang et al., 2022; 2023; Villegas et al., 2022). Our architecture consists of a pre-trained text-conditioned masked motion model and a motion control model. The pre-trained model provides a strong motion prior based on text prompts, while the motion control model introduces additional spatial control signals. Specifically, the motion control model is a trainable replica of the pre-trained masked motion model, as shown in Fig 3. Each Transformer layer in the original model is paired with a corresponding layer in the trainable copy, connected via a zero-initialized linear layer. This initialization ensures that the layers have no effect at the start of training. Unlike the original masked motion model, the motion control model incorporates two conditions: the text prompt $W$ from the pre-trained CLIP model (Radford et al., 2021) and the spatial control signal $S$. The text prompt $W$ influences the motion tokens through attention, while the spatial signal $S$ is directly added to the motion token sequence via a projection layer.
@@
Generative Masking Training with Consistency Feedback. The conditioned masked transformer is trained to learn the conditional distribution $p_{\theta}\left(x_{i}\mid X_{\overline{\mathbf{M}}},W,S\right)$ by reconstructing the masked motion tokens, conditioned on the unmasked tokens $X_{\overline{\mathbf{M}}}$, text prompt (W), and spatial control signal (S). The spatial control condition is a sequence of joint control signals $S=[s_{1},s_{2},...,s_{T}]$ with $s_{i}\in\mathbb{R}^{22\times 3}$. Each control signal $s_{i}$ specifies the targeted 3D coordinates of the joints to be controlled, among the total 22 joints, while joints that are not controlled are zeroed out. Since the semantics of the generated motion are primarily influenced by the textual description, to guarantee the controllability of spatial signals, we extract the spatial control signals from the generated motion sequence and directly optimize the consistency loss between input control signals and those extracted from the output. This consistency training not only enhances controllability but also addresses a unique challenge in controllable motion generation. In the image domain, spatial control signals can be directly applied, and uncontrolled regions are simply zeroed out. However, for motion control, zero-valued 3D joint coordinates may either indicate that a joint should be controlled, with its target position at the origin in Euclidean space, or that the joint is uncontrolled. Consistency training explicitly guides the model on which joints are controlled and which are not, effectively resolving this ambiguity.
Training-time Differential Sampling. While consistency training offers significant benefits, integrating consistency loss into the training of generative masked models presents a challenge: the need to convert motion tokens in the latent space into motion representations in Euclidean space. This conversion requires sampling from the categorical distribution of motion tokens during training, a process that is inherently non-differentiable. To address this, we leverage the straight-through Gumbel-Softmax technique (Jang et al., 2017). This approach performs categorical sampling during the forward pass and approximates the categorical distribution with differentiable sampling using the continuous Gumbel-Softmax distribution during the backward pass, i.e.,
$p_{\theta}\left(x_{i}\mid X_{\overline{\mathbf{M}}},W,S\right)=\frac{\exp\left((\ell_{i}+g_{i})/\tau\right)}{\sum_{j=1}^{k}\exp\left(\ell_{j}/\tau\right)},$ (3)
where $\tau$ refers to temperature and $g$ represents Gumbel noise with $g_{1},\dots,g_{k}$ being independent and identically distributed (i.i.d.) samples from a $\text{Gumbel}(0,1)$ distribution. The $\text{Gumbel}(0,1)$ distribution can be sampled via inverse transform sampling by first drawing $u\sim\text{Uniform}(0,1)$ and then computing $g=-\log(-\log(u))$.
Motion Consistency Loss. With the help of the training-time differential sampling, we are able to define the consistency loss, which assesses how closely the joint control signal extracted motion the generated motion aligns with the input spatial control signal $s$:
@@
where $\sigma_{nj}$ is a binary value indicating whether the spatial control signal $s$ contains a control value at frame $n$ for joint $j$. The motion tokenizer decoder $D(\cdot)$ converts motion embedding into relative position in local coordinate system and $R(\cdot)$ further transforms the joint’s local positions to global absolute locations. $e_{c}$ are the entry embeddings of the codebook and $c$ are the entry indices that are sampled from the motion distribution $p_{\theta}\left(c\mid X_{\overline{\mathbf{M}}},W,S\right)$. The global location of the pelvis at a specific frame can be calculated from the cumulative aggregation of rotations and translations from all previous frames. The locations of the other joints can also be computed by the aggregation of the relative positions of the other joints to the pelvis position. The final loss for masked consistency training is the weighted combination masked training loss and motion consistency loss:
$\mathcal{L}=\alpha\mathcal{L}_{\text{mask}}+(1-\alpha)L_{s}(e_{c},s).$ (5)
3.3 Inference-time Logits and Codebook Editing
Figure 4: Inference of ControlMM. Spatial control is added to the model as input. The output logits are reconstructed and optimized through Differentiable Sampling in each iteration.
To achieve more accurate and generalizable spatial control, we propose directly optimizing the logits of the motion token classifier along with the motion codebook, while keeping the rest of the network frozen. The goal is to enhance control precision by further reducing the discrepancy between the generated motion and the desired control objectives. This approach does not require pretraining on specific spatial control signals, allowing the model to handle arbitrary, out-of-distribution spatial signals during inference, enabling new control tasks such as obstacle avoidance in a zero-shot manner.
The core idea behind logit editing is to update the learned logits through gradient-guided optimization during inference, allowing manipulation of the conditional motion distribution. This ensures that the generated motion, sampled from the adjusted distribution, aligns closely with the input control signals. The optimization process is initialized with the logits obtained from masked consistency training, and these logits are iteratively updated to minimize the consistency loss.

### v2

# Spatio-Temporal Control for Masked Motion Synthesis

@@ Abstract
Recent advances in motion diffusion models have enabled spatially controllable text-to-motion generation. However, these models struggle to achieve high-precision control while maintaining high-quality motion generation. To address these challenges, we propose MaskControl, the first approach to introduce controllability to the generative masked motion model. Our approach introduces two key innovations. First, Logits Regularizer implicitly perturbs logits at training time to align the distribution of motion tokens with the controlled joint positions, while regularizing the categorical token prediction to ensure high-fidelity generation. Second, Logit Optimization explicitly optimizes the predicted logits during inference time, directly reshaping the token distribution that forces the generated motion to accurately align with the controlled joint positions. Moreover, we introduce Differentiable Expectation Sampling (DES) to combat the non-differential distribution sampling process encountered by logits regularizer and optimization. Extensive experiments demonstrate that MaskControl outperforms state-of-the-art methods, achieving superior motion quality (FID decreases by  77%) and higher control precision (average error 0.91 vs. 1.08). Additionally, MaskControl enables diverse applications, including any-joint-any-frame control, body-part timeline control, and zero-shot objective control. Video visualization can be found at https://www.ekkasit.com/ControlMM-page/
@@ 1 Introduction
Text-driven human motion generation has recently gained significant attention due to the semantic richness and intuitive nature of natural language descriptions. This approach has broad applications in animation, film, virtual/augmented reality (VR/AR), and robotics. While text descriptions offer a wealth of semantic guidance for motion generation, they often fall short in providing precise joint control over specific human joints, such as the pelvis and hands. As a result, achieving natural interaction with the environment and fluid navigation through 3D space remains a challenge.
@@
To tackle this challenge, a few controllable motion generation models have been developed recently to synthesize realistic human movements that align with both text prompts and joint control signals [43, 42, 55, 51]. However, existing solutions face significant difficulties in generating high-fidelity motion with precise and flexible joint control. In particular, current models struggle to support both sparse and dense joint control signals simultaneously. For instance, some models excel at generating natural human movements that traverse sparse waypoints [21, 42], while others are more effective at synthesizing motions that follow detailed trajectories specifying human positions at each time point [51]. Recent attempts to support both sparse and dense joint inputs encounter issues with control precision; the generated motion often is not aligned well with the control conditions [55]. Besides unsatisfied joint flexibility and accuracy, the quality of motion generation in controllable models remains suboptimal, as evidenced by much worse FID scores compared to models that rely solely on text inputs. Moreover, most current methods utilize motion-space diffusion models, applying diffusion processes directly to raw motion sequences. While this design facilitates the incorporation of joint control signals, the redundancy in raw data introduces computational overhead, resulting in slower motion generation speeds.
Current controllable motion generation methods mostly rely on diffusion model [55, 7, 21, 55]. However, these methods exhibit lower control precision and less generation quality. The current SOTA method, TLControl, [51] achieves high-precision control through test-time optimization with a simple feedforward transformer, but this comes at the cost of lower generation quality, as shown in Tab. 1. Moreover, these methods cannot adapt to arbitrary objective functions at inference time in a zero-shot manner.
To address these challenges, we introduce the first method that integrates joint control into generative masked motion models, enabling simultaneous high-fidelity motion generation and precise control across a broad range of tasks, as illustrated in Fig. LABEL:fig:landing. In contrast to diffusion-based approaches, masked models generate motion sequences by training a multi-category token classifier and subsequently sampling from the learned categorical distributions conditioned on input signals, such as text and joint trajectories. Building upon this insight, we propose MaskControl—a straightforward yet powerful control mechanism for generative masked motion models—that implicitly and explicitly manipulates the logits of the token classifier to align token distributions closely with input control signals.
Our contributions can be summarized as follows.
  • We propose MaskControl, the first approach to introduce controllability to generative masked motion models.
  • We introduce two novel control components of MaskControl: (1) Logits Regularizer that implicitly perturbs logits at training time to align the distribution of motion tokens with the controlled joint positions. (2) Logits Optimization explicitly optimizes the predicted logits during inference time, directly reshaping the token distribution to minimize the residual errors between the generated motion and the target joint position. Moreover, we propose Differentiable Expectation Sampling (DES) to overcome the challenge of non-differentiable probabilistic token selection in logits regularization and optimization.
  • We show that Logits Optimization can be generalized to solve unseen control tasks in a zero-shot manner.
@@
  • We conduct extensive qualitative and quantitative evaluations on multiple tasks, demonstrating that our approach outperforms current SOTA in both motion generation quality and control precision while supporting multiple applications i.e. any-joint-any-frame control, body part timeline control, and zero-shot objective control.
Table 1: Comparison of text-conditioned motion generation with joint control signals. Our MaskControl SOTA achieves performance by leveraging Masked Motion Model, demonstrating high-precision control (low Average Error) while maintaining high generation quality (low FID). Previous SOTA methods utilize diffusion in motion space, latent space, or simple feed-forward models. ‘’ indicates the ability to control motion using zero-shot objective functions, while ‘✗’ denotes the lack of this capability. ‘-’ signifies control limited to the pelvis.
Method FID $\downarrow$ Average Error (cm) $\downarrow$ Base Model Zero-shot Objective
GMD [21] 0.576 14.39 Motion Diffusion -
OmniControl [55] 0.218 3.38 Motion Diffusion
MotionLCM [7] 0.531 18.97 Latent Diffusion
TLControl [51] 0.271 1.08 Feed Forward
Ours 0.061 0.98 Masked Model
@@ 2 Related Work
Text-driven Motion Generation. Early methods for text-to-motion generation primarily focus on aligning the latent distributions of motion and language, typically by employing loss functions such as Kullback-Leibler (KL) divergence and contrastive losses. Representative works in this domain include Language2Pose [1], TEMOS [37], T2M [15], MotionCLIP [47], and DropTriple [56]. However, the inherent discrepancy between the distribution of text and motion often results in suboptimal generation quality when using these latent space alignment techniques.
@@
Recently, diffusion models have become a widespread choice for text-to-motion generation, operating directly in the motion space [48, 60, 23], VAE latent space [5], or quantized space [31, 24]. In these works, the model gradually denoises the whole motion sequence to generate the output in the reverse diffusion process. Another line of work explores the token-based models in the human motion domain, for example, autoregressive GPTs [14, 58, 20, 61] and masked motion modeling [40, 39, 16]. These methods learn to generate discrete motion token sequences that are obtained from a pretrained motion VQVAE [12, 53]. While GPT models usually predict the next token from history tokens, masked motion models utilize the bidirectional context to decode the masked motion tokens. By predicting multiple tokens at once, the masked modeling methods can generate motion sequences in as few as 10 steps, achieving state-of-the-art performance on generation quality and efficiency. Despite the performance gains of masked motion models, supporting spatial controllability in these models remains unexploited. This paper is the first work that proposes controllable masked motion model to simultaneously achieve high-quality motion generation with high-precision spatial control.
@@
Controllable Motion Synthesis. In addition to text prompts, synthesizing motion based on other control signals has also been a topic of interest. Example control modalities include music [29, 27, 26, 44, 45, 49], interacting object [25, 9, 28, 2], tracking sensors [11], scene [17, 52] programmable motion [30], style [62], goal-reaching task [10], and multi-track timeline control [38]. [35, 36, 54, 57, 32, 33, 46] incorporate physics into motion generation. To control the trajectory, PriorMDM [43] finetunes MDM to enable control over the locations of end effectors. CondMDI [6] generates motion in-betweening from arbitrarily placed dense or sparse keyframes. GMD [21] and Trace and Pace [42] incorporates spatial control into the diffusion process by guiding the root joint location. OmniControl [55] extends the control framework to any joint, while MotionLCM [7] applies this control in the latent space, both leveraging ControlNet [59]. DNO [22] introduces an optimization process on the diffusion noise to generate motion that minimizes a differentiable objective function. Recent approaches [51, 18] model each body part separately to achieve fine-grained control but are limited to dense trajectory objectives. The existing frameworks for controllable motion generation predominantly rely on diffusion models; however, these models typically suffer from low-quality motion generation.
Figure 1: Overall architecture of MaskControl. (a) Motion Tokenizer transforms the motion sequence into discrete motion tokens. (b) Differentiable Expectation Sampling (DES) is a differentiable sampling from logits enabling differentiable conversion between discrete tokens in codebook space and transformer token space. (c) Training: Logits Regularizer ensures high-quality motion by generating embedding closely aligns with joint control signals during an unmasking process. (d) Inference: Logits Optimization guides logits during the unmasking process at inference time based on the objective function.
3 MaskControl
@@
The objective of MaskControl is to enable controllable text-to-motion generation based on a masked motion model that generates high-precision and high-quality motion. In particular, given a text prompt and an additional joint control signal, our goal is to generate a physically plausible human motion sequence that closely aligns with the textual descriptions, while following the joint control conditions, i.e., $(x,y,z)$ positions of each human joint at each frame in the motion sequence. Towards this goal, in Sec. 3.1, we first introduce the background of conditional motion synthesis based on the generative masked motion model. We then describe two key components of MaskControl, including Logits Regularizer in Sec. 3.2 and inference-time Logits Optimization in Sec. 3.3. The first component aims to learn the categorical distribution of motion tokens, conditioned on joint control signals during training time. The second component aims to improve control precision by optimally modifying learned motion distribution via Logits Optimization during inference time. Lastly, we introduce Differentiable Expectation Sampling to overcome non-differentiability of categorical sampling during Logits Regularizer and Optimization.
3.1 Preliminary: Masked Motion Model
@@
Generative masked motion models  [40, 39, 16] generally consist of two stages: Motion Tokenizer and Text-conditioned Masked Transformer. The objective of the Motion Tokenizer is to learn a discrete representation of motion by quantizing the encoder’s output embedding $z$ into a codebook $\mathcal{C}$. For a given motion sequence $\mathcal{P}=[p_{1},p_{2},...,p_{F}]$, where each frame $p$ represents a 3D pose, Motion Tokenizer outputs a discrete motion tokens $X=[x_{1},x_{2},...,x_{L}]$. Specifically, the encoder compresses $\mathcal{P}$ into a latent embedding $z\in\mathbb{R}^{t\times d}$ with a downsampling rate of $F/L$. The embedding $z$ is quantized into codes $c\in\mathcal{C}$ from the codebook $\mathcal{C}=\{c_{k}\}_{k=1}^{K}$, which contains $K$ codes. The nearest code is selected by minimizing the Euclidean distance between $z$ and the codebook entries, computed as $\hat{z}=\operatorname{argmin}_{j}\left\|\mathbf{z}-c_{j}\right\|_{2}^{2}$. The vector quantization loss $L_{VQ}$ is defined as:
$L_{VQ}=\|\operatorname{sg}(\mathbf{z})-\mathbf{c_{j}}\|_{2}^{2}+\beta\|\mathbf{z}-\operatorname{sg}(\mathbf{c_{j}})\|_{2}^{2},$ (1)
@@
During the second stage, the quantized motion token sequence $X=[x_{1},x_{2},...,x_{L}]$ is updated with [MASK] tokens to form the corrupted motion sequence $X_{\overline{\mathbf{M}}}$. This corrupted sequence along with text embedding $W$ are fed into a text-conditioned masked transformer parameterized by $\theta$ to reconstruct input motion token sequence with reconstruction probability equal to $p_{\theta}\left(x_{i}\mid X_{\overline{\mathbf{M}}},W\right)$, which is obtained by the motion token classifier. The objective is to minimize the negative log-likelihood of the predicted masked tokens conditioned on text:
@@
$\small\mathcal{L}_{\text{mask}}=-\underset{\mathbf{X}\in\mathcal{D}}{\mathbb{E}}\left[\sum_{\forall i\in[1,L]}\log p\left(x_{i}\mid X_{\overline{\mathbf{M}}},W\right)\right].$ (2)
@@
During inference, the transformer masks out the tokens with the least confidence and re-samples these tokens according to their respective distributions $p_{\theta}\left(x_{i}\mid X_{\overline{\mathbf{M}}},W\right)$in the subsequent step. The number of masked tokens $n_{M}$ is controlled by a masking schedule, a decaying function of the step $t$. Early steps use a large masking ratio due to high uncertainty, and as the process continues, the ratio decreases as more context is available from previous predictions.
3.2 Logits Regularizer
@@
MaskControl aims to generate a human motion sequence based on the text prompt $W$ and joint control signals $S$. Towards this goal, we introduce Logits Regularizer to conditioned masked transformer, which aims to learn the motion token distribution jointly conditioned both on $W$ and $S$. Logits Regularizer implicitly alters the output logits of pre-trained text-to-motion model, thus changing the distribution of motion tokens, toward joint control positions.
@@
Model Architecture. Diffusion ControlNet has shown its excellence in adding additional control signals to pretrained image diffusion model [59]. We demonstrate for the first time that similar design principle can be applied to introduce controllability to generative masked models [8, 3, 4, 50]. Our architecture consists of a pre-trained text-conditioned masked motion model and a Logits Regularizer. The pre-trained model provides a strong motion prior based on text prompts, while the Logits Regularizer introduces additional joint control signals. Specifically, the Logits Regularizer is a trainable replica of the pre-trained masked motion model, as shown in Fig 1. Each Transformer layer in the original model is paired with a corresponding layer in the trainable copy, connected via a zero-initialized linear layer. This initialization ensures that the layers have no effect at the start of training. Unlike the original masked motion model, the Logits Regularizer incorporates two conditions: the text prompt $W$ from the pre-trained CLIP model [41] and the joint control signals $S$. The text prompt $W$ influences the motion tokens through attention, while the joint control signal $S$ is directly added to the motion token sequence via a projection layer.
@@
Joint Control Signal. The conditioned masked transformer is trained to learn the conditional distribution $p_{\theta}\left(x_{k}\mid X,X_{\overline{\mathbf{M}}},W,S\right)$ by reconstructing the masked motion tokens $X_{\overline{\mathbf{M}}}$, conditioned on the unmasked tokens $X$, text prompt W, and joint control signals S. The joint control condition is a sequence of joint control signals $S=[s_{1},s_{2},...,s_{F}]$ with $s_{i}\in\mathbb{R}^{j\times 3}$. Each control signal $s_{i}$ specifies the targeted 3D coordinates of the joints to be controlled, among the total $j$ joints, while joints that are not controlled are zeroed out. Since the semantics of the generated motion are primarily influenced by the textual description, to guarantee the controllability of joint control signals, we extract the joint control signals from the generated motion sequence and directly optimize the consistency loss between input control signals and those extracted from the output. Note that zero-valued in 3D joint coordinates can be ambiguous. Please refer to Sec. A.11 for more details.
Motion Consistency Loss evaluates the alignment between the generated motion and the input joint control signals $s$:
@@
where $\sigma_{nj}$ is a binary value indicating whether the joint control signals $s$ contains a control value at frame $n$ for joint $j$. The motion tokenizer decoder $D(\cdot)$ converts motion embedding into relative position in local coordinate system and $R(\cdot)$ further transforms the joint’s local positions to global absolute locations. The global location of the pelvis at a specific frame can be calculated from the cumulative aggregation of rotations and translations from all previous frames. The locations of the other joints can also be computed by the aggregation of the relative positions of the other joints to the pelvis position.
Logits Consistency Loss extends the objective function of masked motion model in Eq. 2 by applying negative log-likelihood to all positions, including unmasked ones, conditioned on both the text $W$ and joint control signals $S$.
$\small\mathcal{L}_{\text{logits}}=-\sum_{\forall i\in[1,L]}\log p\left(x_{i}\mid X_{\overline{\mathbf{M}}},W,S\right).$ (4)
The final loss is the weighted combination Logits Consistency Loss and Motion Consistency Loss:
$\mathcal{L}=\alpha\mathcal{L}_{\text{logits}}+(1-\alpha)L_{s}(e_{c},s).$ (5)
3.3 Logits Optimization
The goal of inference-time Logits Optimization is to enhance control precision by further reducing the discrepancy between the generated motion and the desired control objectives. This approach does not require pretraining on specific tasks, allowing the model to handle arbitrary objective functions during inference, enabling new control tasks in a zero-shot manner.
The core idea behind Logits Optimization is to update the learned logits through gradient-guided optimization during inference, perturbing the motion token distribution. Optimizing logits makes optimization during unmasking process possible since the process requires conversion between transformer token for unmask sampling and codebook embedding for motion reconstruction, as shown in Fig. 1. The optimization process is initialized with the logits obtained from conditioned masked transformer with Logits Regularizer, and these logits are iteratively updated to minimize the motion consistency loss $L_{s}$.